# Adaptive Reduced Multilevel Splitting

### v1

@@ 1.3 Sketch of the proposed method
  • The simulation of importance or proposal distributions $\mu_{l}$ at various level $l$ is implemented using an AMS Monte Carlo methodology, in which $\mu_{l}$ is approximated by the empirical distribution of a sample of particles.
  • At each iteration, the snapshot $X^{(k+1)}$ is sampled among the sample of particles (which become candidates) distributed according to the importance distribution $\mu_{l^{(k)}}$, with a well-chosen level $l^{(k)}$. It will be necessary in early stages of the algorithm to replace the uniform sampling of $X^{(k+1)}$ among particles by a penalized sampling based on a learning function, for instance by adding probability weights to states with high errors, as discussed in Section 2.3.
  • The true model is evaluated at $X^{(k+1)}$ and the reduced model is updated accordingly.
  • Importance sampling estimators of the target rare event problem are provided using the sequence of snapshots $(X^{(k)},\Psi^{\ast}(X^{(k)}))_{k\geq 1}$.
The main methodological contribution of the present work is the specific choice of the level $l^{(k)}$. At each iteration $k$, we will use the information given by the error quantification in order to adaptively choose a level $l^{(k)}$ so that $\mu_{l^{(k)}}^{(k)}$ remains a reasonable importance distribution associated with the true target $\eta^{\ast}_{l^{(k)}}$. This way, we follow (using the AMS methodology here) the concept of rare event sequential sampling along the flow of targets $l\mapsto\eta^{\ast}_{l}$. However, because we rely on a surrogate model when sampling according to $\mu_{l^{(k)}}^{(k)}$, we stop the increase of rareness of the targets when the reduced modeling becomes critically inaccurate, the associated level being denoted $l:=l^{(k)}$. Snapshot (true model) sampling is then preformed using a sample of candidates following $\mu_{l^{(k)}}^{(k)}$, considered at the precise level $l^{(k)}$ of the flow where increasing the quality of the reduced model becomes necessary. By doing so, we ensure that the choice of the snapshot will be consistent with the unknown target, albeit for a moderately large level $l^{(k)}$.
The evaluation of the accuracy of an importance distribution will be done using a quantification of the following logarithmic cost of importance sampling by relative entropy:
$\ln\mathrm{cost}\simeq\mathrm{Ent}(\eta_{l^{(k)}}\mid\mu^{(k)}_{l^{(k)}})\displaystyle\mathop{=}^{{\rm def}}\eta_{l^{(k)}}(\ln\frac{\eta_{l^{(k)}}}{\mu^{(k)}_{l^{(k)}}}),$
@@
discussed and justified in Appendix A. The latter cannot be evaluated exactly, but we can leverage the existence of error estimates in order to obtain (more or less pessimistic) approximations of it.
Remark that the more one learns about the true score $S^{\ast}$ by generating snapshots, the more we improve the reduced model (for specific, well-chosen rare states), the more we may reduce the cost of importance sampling. When this happens, the log-cost criteria limiting the level $l$ to $l^{(k)}$ is alleviated, and the algorithm is likely to increase the level beyond $l^{(k)}$ in subsequent iterations.
@@
The proposed method enables to obtain importance sampling estimators. If for instance the snapshots are sampled according to importance distribution $X^{k+1}\sim\mu^{(k)}_{l^{(k)}}$ from the start, it naturally leads to the following estimator:
@@
which by construction, is an unbiased empirical estimator of the reference distribution $\pi$ (see Section 3.2), up to sampling errors of the proposal $\mu_{l}$ by AMS. This leads to unnormalized estimators of the target $\eta^{\ast}_{l}$ by evaluating
$\hat{\pi}^{(k)}(\varphi\,{\rm l}{\rm l}_{S^{\ast}>l}),$
for various $l$ and test function $\varphi$, which can be easily done because the true score of each snaphots has been evaluated preliminarily by construction of the algorithm and in particular
@@
As mentioned above, the sequence of importance distribution $\mu_{l^{(k)}}^{(k)}$ will be simulated by a population of particles and an AMS methodology in the spirit of current rare event simulation algorithms. The only difference is that the target distribution is constructed with a reduced score that is adaptively enriched by few evaluations of the true score throughout the simulation. An important remark is that updating the reduced score will in general break the nesting of the sequence of distributions. A simple way of addressing this problem is to perform a fresh AMS simulation each time that a new snapshot is picked and that the reduced score is updated. To significantly reduce the number of evaluations of the reduced model, another option is to propose a bridging procedure. This is the second main methological contribution of this work. At iteration $k$, this procedure searches in the simulation history for a well-chosen previous proposal denoted $\mu_{l^{(k_{\mathrm{bridge}})}}^{(k_{\mathrm{bridge}})}$ with $k_{\mathrm{bridge}}\leq k$, so that the updated proposal denoted $\mu_{l_{\mathrm{bridge}}(k+1)}^{(k+1)}$, with ${l_{\mathrm{bridge}}(k+1)}$ a well-designed level, is dominated by and sufficiently close to the former. The latter proposal is then simulated starting from the former one using an AMS framework.
In numerical experiments, we study the proposed algorithm, called Adaptive Reduced Multilevel Splitting (ARMS) in the context of these two variants: fresh AMS simulation or bridging. We show the empirical convergence of the ARMS algorithm on a toy model and a rare-event problem based on a more realistic PDE. Our study confirms that the proposed relative entropy criterion plays a crucial role in ARMS and prevent the simulation to be taken into spurious regions of the state-space remote from the region of interest, resulting in a large variance of the IS estimator. Moreover, by evaluating the computational complexity in relation to the squared error of the estimate, we show empirically that the bridging procedure significantly reduces the simulation cost while achieving a comparable squared error.
The paper is organized as follows. The main methodological concepts underlying the ARMS algorithm are presented in Section 2. We then introduce an idealized version of the ARMS algorithm in Section 3, which assumes the exact evaluation of expectations and exact sampling of distributions. In Section 4, we define ARMS, which consists in a practical implementation of this idealized algorithm using AMS simulations. In Section 5, we numerically evaluate the proposed algorithm for two rare event problems. A final section deals with conclusions and perspectives. The appendix collects some technical proofs and details.
@@ 2.1 Importance sampling cost
The domination requirement, denoted
$\mu_{l}\gg\eta_{l}^{\ast},$
holds true by definition if and only if one has the inclusion $\left\{S^{\ast}>l\right\}\subset\left\{S>l\right\}$. This is not guaranteed here as $S$ is an approximation of the true score $S^{\ast}$, and since $S^{\ast}$ is not computed for each sample states, one cannot check it in a systematic fashion. However, as we will discuss in the next section, the error bound $E$ can be used to propose a sufficient condition corresponding to a worst case scenario.
Unfortunately, even if the above domination condition holds, importance sampling becomes quickly infeasible if the target distribution diverges too widely from the proposal, a typical phenomenon in high dimension. As discussed in Appendix A, it has been thoroughly argued in [7] (see also [6]) that the cost of importance sampling – in terms of the required sample size $K$– is quite generically and roughly given by the exponential of the relative entropy between $\eta^{\ast}_{l}$ and $\mu_{l}$:
@@
Note that domination between measures is implied by finite relative entropy, and will rather use throughout the paper the condition
$\mathrm{Ent}(\eta\mid\mu)<+\infty$
in order to check the domination condition $\mu\gg\eta$.
@@ 2.2 First non-achievable level
In this section, we describe how to choose the target level $l^{(k)}$, which is suited to the current score function $S^{(k)}$ at iteration $k$ ($k$ being the number of newly sampled snapshots of the true model) of the main loop of the algorithm. The idea is to tune $l^{(k)}$ with a criteria evaluating that surpassing this level should require, in some sense, a better surrogate model, and thus new snapshots. The adjusted proposal $\mu^{(k)}_{l^{(k)}}$ will then be used to a sample a new snapshot, and obtain an updated score function $S^{(k+1)}$ for the next iteration $k+1$. In short, we are going to develop a heuristic criterion on levels to detect when a new snapshot is needed.
@@
The general insight is the following. Consider the following simpler problem: one is given a family of rare event target distributions $\left(\eta^{\ast}_{l}\right)_{l\in\mathbb{R}}$ in the form of (1.1), a given computational budget, and a family of importance sampling proposal distributions $(\mu_{l}^{(k)})_{l\in\mathbb{R}}$ in the form of (1.3). One wants to perform some sequential algorithm to sample the distributions $\mu_{l}$ along an increasing sequence of levels $l$ as high as possible, and then use importance sampling. However, we do not want to waste resources for levels that are ”unreasonably” high. As an idealized criteria giving achievable levels, we propose the importance sampling cost (2.1) of the true target. The first non-achievable (hence critical) level is thus given when the latter matches a certain given numerical threshold $e$: loosely speaking,
$\text{Find first}\,l:\quad\mathrm{Ent}(\eta^{\ast}_{l}\mid\mu^{(k)}_{l})\geq e.$
@@
By doing so, we ensure that the snapshot sampled with respect to $\mu_{l}^{(k)}$ will not diverge too much from a meaningful target distribution $\eta^{\ast}_{l}$ with an appropriate level $l$. As an example, if, as it is the case in the present context, $\mu_{l}^{(k)}$ has been constructed using a reduced model based on $k$ snapshots $(X_{1},\ldots,X_{k})$, it is unlikely that it can be used as an importance proposal for levels higher than $\max(S^{\ast}(X_{1}),\ldots,S^{\ast}(X_{k}))$, as the quality of the approximation $S^{\ast}-S^{(k)}$ will deteriorate in areas with higher scores.
@@
In a practical context, the log cost $\mathrm{Ent}(\eta^{\ast}_{l}\mid\mu_{l}^{(k)})$ has to be estimated using some prior information on $\eta^{\ast}_{l}$, for instance using some form of error quantification. We propose in this work the following method, which is by no means unique. We consider a pessimistic (respectively an optimistic) surrogate of $\eta^{\ast}_{l}$, denoted by $\check{\eta}_{l}^{(k)}$ (respectively by $\hat{\eta}_{l}^{(k)}$), constructed with the reduced score $S^{(k)}$ and its error quantification $E^{(k)}$, and typically associated with a much lower (respectively higher) rare event probability than $\eta^{\ast}_{l}$. In the present work, we will choose (2.3)-(2.4) which even satisfy:
$\check{\eta}_{l}^{(k)}\ll\eta^{\ast}_{l}\ll\hat{\eta}_{l}^{(k)}$
@@
The first non-achievable level will be estimated using the pessimistic surrogate by searching the closest level such that the log cost $\mathrm{Ent}\left(\check{\eta}_{l}^{(k)}\mid\mu_{l}^{(k)}\right)$ matches the threshold $e$. Given an initial level $l_{0}$, this yields the following rigorous definition of the (critical) first non-achievable level:
$l^{(k)}(e,l_{0})=\min\{l\geq l_{0}:\;\mathrm{Ent}\left(\check{\eta}_{l}^{(k)}\mid\mu_{l}^{(k)}\right)\geq e\;\;\text{or}\;\;\mathrm{Ent}\left(\hat{\eta}_{l_{\mathrm{max}}}^{(k)}\mid\mu_{l}^{(k)}\right)=+\infty\}.$ (2.2)
@@
$l(e,l_{0})$ is a level such that, in the (unfortunate) case where the true target is given by the pessimistic choice $\check{\eta}_{l}^{(k)}$, a log cost of $e$ is necessary and sufficient in order to perform importance sampling.
Note that the last inequality constraint in (2.2) ensures that the support of $\mu_{l}^{(k)}$ includes the support of the final target ${\eta}^{\ast}_{l_{\mathrm{max}}}$. A more conservative and rigorous variant consists in requiring
$\mathrm{Ent}\left(\hat{\eta}_{l}^{(k)}\mid\mu_{l}^{(k)}\right)<+\infty$
instead of $l_{\mathrm{max}}$. The latter more conservative condition rigorously implies that $\eta^{\ast}_{l}\ll\mu_{l}^{(k)}$, and in the case where $\check{\eta}_{l}^{(k)},\hat{\eta}_{l}^{(k)}$ are defined by indicator densities, $l(e,l_{0})$ is then exactly the unique level that set a quantitative guarantee on the worst-case log cost of importance sampling the true target $\eta^{\ast}_{l}$ (the largest cost in a pessimistic scenario given by the error bound). We have relaxed this condition because, especially early in the algorithm, we are not interested in artificial constraints on sampling areas of the state space that we know will become irrelevant later for larger $l$.
@@
In the present rare event case where for $l\in\mathbb{R}$, the target is given by (1.1) and the importance distribution by (1.3) with pointwise error bounds (1.2), we propose the pessimistic and optimistic surrogate targets
@@
Using the surrogate (2.3), the particularization of the worst-case log cost in (2.2) is
$\mathrm{Ent}\left(\check{\eta}_{l}^{(k)}\mid\mu_{l}^{(k)}\right)=\ln\frac{Z_{l}^{(k)}}{\check{Z}_{l}^{(k)}},$
thus the logarithm of the ratio of the approximate probability ${\pi\left(\left\{S^{(k)}>{l}\right\}\right)}$ over the pessimistic probability ${\pi\left(\left\{S^{(k)}-E^{(k)}>{l}\right\}\right)}$.
@@
It can easily be checked that the critical level $l^{(k)}(e,l_{0})$ tends to infinity when the error $E^{(k)}$ vanishes pointwise.
In summary, at iteration $k$, starting from some level $l_{0}$, the target level $l^{(k)}$ is chosen such that the log cost for importance sampling of a more pessimistic surrogate target $\check{\eta}_{l^{(k)}}$ is given by a prescribed budget $e$, i.e.,
$l^{(k)}=l^{(k)}(e,l_{0}),$
as defined by by (2.2). We will consider different settings for $l_{0}$ in the full algorithm as explained later in Section 4.3.

### v3

@@ 1.3 Contributions
The Monte Carlo aspect of the algorithm is implemented using an AMS simulation methodology, in which $\mu_{l}^{(k)}$ is approximated by the empirical distribution of a sample (population) of particles (clones).
Our first main methodological contribution is the specific criterion used to adaptively compute the maximum level of iteration $k$ denoted $l^{(k)}$. At each iteration $k$, we will use the information given by the error quantification (1.2) so that $\mu_{l^{(k)}}^{(k)}$ remains a reasonable importance distribution for sampling the true target $\eta^{\ast}_{l^{(k)}}$. The evaluation of the accuracy of an importance distribution will be done using a quantification of the logarithmic cost of importance sampling by relative entropy:
$\ln\mathrm{cost}\simeq\mathrm{Ent}(\eta_{l^{(k)}}\mid\mu^{(k)}_{l^{(k)}})\displaystyle\mathop{=}^{{\rm def}}\eta_{l^{(k)}}(\ln\frac{\eta_{l^{(k)}}}{\mu^{(k)}_{l^{(k)}}}).$
@@
This choice is discussed and justified in Appendix A. The latter cannot be evaluated exactly, but we can leverage the existence of error estimates in order to obtain (more or less pessimistic) approximations of it. Relying on this cost, we are able to stop the increase of rareness of the targets when the reduced modeling becomes critically inaccurate, the associated maximum level being $l^{(k)}$.
Snapshot sampling is then preformed using the distribution $\mu_{l^{(k)}}^{(k)}$, which we may reweight by favoring states with higher error quantification. Remark that snapshot sampling is performed at the precise level $l^{(k)}$ of the flow of reduced proposal distributions $(\mu_{l}^{(k)})_{l\in\mathbb{R}}$ where increasing the quality of these proposals becomes necessary: this prevents a poor sampling of a new snapshot inconsistent with the unknown target while keeping the associated level $l^{(k)}$ as large as possible.
@@
The proposed method can also yield importance sampling estimators. If for instance the snapshots are sampled from the start without re-weighting according to proposal $X^{k+1}\sim\mu^{(k)}_{l^{(k)}}$, it naturally leads to the following estimator:
@@
which by construction, is an unbiased empirical estimator of the reference distribution $\pi$ (see Section 2.6 for a proof), up to sampling errors of the proposal $\mu_{l}$, which will be done in practice using AMS. This leads to an IS estimator e.g. of the rare event probability
@@
As mentioned above, the sequence of importance distribution $\mu_{l^{(k)}}^{(k)}$ will be simulated by a population of particles and an AMS methodology in the spirit of current rare event simulation algorithms. The only difference is that the target distribution is constructed with a reduced score that is adaptively enriched by few evaluations of the true score throughout the simulation. As remarked in Section 1.2, updating the reduced score will in general break the nesting of the sequence of distributions. A simple way of addressing this problem is to perform a fresh AMS simulation each time that a new snapshot is picked and that the reduced score is updated. To significantly reduce the number of evaluations of the reduced model, another option is to propose a bridging procedure. This is the second main methodological contribution of this work. At iteration $k$, this procedure searches in the simulation history for a well-chosen previous proposal denoted $\mu_{l^{(k_{b})}}^{(k_{b})}$ with $k_{b}\leq k$, so that the updated proposal denoted $\mu_{l_{0}^{(k+1)}}^{(k+1)}$, with ${l_{0}^{(k+1)}}$ a well-designed level, is dominated and sufficiently close to the former. The latter proposal is then simulated starting from the former one using an AMS framework.
In numerical experiments, we study the proposed algorithm, called Adaptive Reduced Multilevel Splitting (ARMS) and show the empirical convergence of the ARMS algorithm on a toy model and a rare-event problem based on a realistic elliptic PDE and a reduced basis approach for model reduction. Our study confirms that the proposed relative entropy criterion plays a crucial role in ARMS and prevent the simulation to be taken into spurious regions of the state-space remote from the region of interest, resulting in a large variance of the IS estimator. Moreover, by evaluating the computational complexity in relation to the squared error of the estimate, we show empirically that the bridging procedure reduces in a certain regime the simulation cost while achieving a comparable squared error.
The paper is organized as follows. The main methodological concepts underlying the ARMS algorithm are presented in Section 2. We end this section by introducing an idealized version of the ARMS algorithm, which assumes the exact evaluation of expectations and exact sampling of distributions. In Section 3, we define ARMS, which consists in a practical implementation of this idealized algorithm using AMS simulations. In Section 4, we numerically evaluate the proposed algorithm for various rare event problems. A final section deals with conclusions and perspectives. The appendix collects some technical proofs and details.
@@ 2.1 Importance sampling cost
The domination requirement, denoted $\mu_{l}\gg\eta_{l}^{\ast},$ holds true by definition if and only if one has the inclusion $\left\{S^{\ast}>l\right\}\subseteq\left\{S>l\right\}$. This is not guaranteed here as $S$ is an approximation of the true score $S^{\ast}$, and since $S^{\ast}$ is not computed for each sample states, one cannot check it in a systematic fashion. Moreover, even if the above domination condition holds, importance sampling becomes quickly infeasible if the target distribution diverges too widely from the proposal, a typical phenomenon in high dimension. As discussed in Appendix A, it has been thoroughly argued in [8] (see also [7]) that the cost of importance sampling – in terms of the required sample size $K$– is quite generically and roughly given by the exponential of the relative entropy between $\eta^{\ast}_{l}$ and $\mu_{l}$:
@@
Note that domination between measures is implied by finite relative entropy, and will rather use throughout the paper the condition $\mathrm{Ent}(\eta\mid\mu)<+\infty$ in order to check the domination condition $\mu\gg\eta$.
@@ 2.2 First non-achievable level
In this section, we describe how to choose the target level $l^{(k)}$, which is suited to the current score function $S^{(k)}$ at iteration $k$ of the algorithm – $k$ being the number of sampled snapshots of the true model. The idea is to tune $l^{(k)}$ with a criteria evaluating that surpassing this level should require, in some sense, a better surrogate model, and thus new snapshots. The adjusted proposal $\mu^{(k)}_{l^{(k)}}$ will then be used to sample a new snapshot, and obtain an updated score function $S^{(k+1)}$ for the next iteration $k+1$. In short, we are going to develop a heuristic criterion on levels to detect when a new snapshot is needed.
@@
The general insight enabling to define $l^{(k)}$ is the following. Consider the following problem: one is given a family of rare event target distributions $\left(\eta^{\ast}_{l}\right)_{l\in\mathbb{R}}$ in the form of (1.1), a given computational budget, and a family of importance sampling proposal distributions $(\mu_{l}^{(k)})_{l\in\mathbb{R}}$ in the form of (1.3). One wants to perform some sequential algorithm to sample the distributions $\mu_{l}$ along an increasing sequence of levels $l$ as high as possible, and then use importance sampling. However, we are given limited resources preventing sampling for levels that are unreasonably high. As an idealized criteria formalizing achievable levels, we propose the importance sampling cost (2.1) of the true target. The (ideal) first non-achievable (hence critical) level is thus given by the smallest level for which this cost matches a certain given numerical threshold $c$, that is:
$\text{Find first}\,l:\quad\mathrm{Ent}(\eta^{\ast}_{l}\mid\mu^{(k)}_{l})\geq c.$
@@
By doing so, we ensure that the snapshot sampled with respect to $\mu_{l^{(k)}}^{(k)}$ will not diverge too much from the target distribution $\eta^{\ast}_{l^{(k)}}$. As an example, if, as it is the case in the present context, $\mu_{l}^{(k)}$ has been constructed using a reduced model based on $k$ snapshots $(X_{1},\ldots,X_{k})$, the first non-achievable level $l^{(k)}$ will typically be of order $\max(S^{\ast}(X_{1}),\ldots,S^{\ast}(X_{k}))$, as the quality of the approximation $S^{\ast}-S^{(k)}$ will deteriorate in areas with higher scores.
@@
In a practical context, the log cost $\mathrm{Ent}(\eta^{\ast}_{l}\mid\mu_{l}^{(k)})$ has to be approximated using some prior information on $\eta^{\ast}_{l}$, for instance using some form of error quantification. We propose in this work the following method, which is by no means unique. We consider a pessimistic (respectively an optimistic) surrogate of $\eta^{\ast}_{l}$, denoted by $\check{\eta}_{l}^{(k)}$ (respectively by $\hat{\eta}_{l}^{(k)}$), constructed with the reduced score $S^{(k)}$ and its error quantification $E^{(k)}$, and typically associated with a much lower (respectively higher) rare event probability than $\eta^{\ast}_{l}$. In the present work, we will choose (2.3)-(2.4) below which satisfy:
$\check{\eta}_{l}^{(k)}\ll\eta^{\ast}_{l}\ll\hat{\eta}_{l}^{(k)}.$
@@
The first non-achievable level will be estimated using the pessimistic surrogate by searching the closest level such that the log cost $\mathrm{Ent}\left(\check{\eta}_{l}^{(k)}\mid\mu_{l}^{(k)}\right)$ matches the threshold $c$. Given an initial level $l_{b}$, this yields the following rigorous definition of the (critical) first non-achievable level:
$l^{(k)}(c,l_{b})=\min\{l\geq l_{b}:\;\mathrm{Ent}\left(\check{\eta}_{l}^{(k)}\mid\mu_{l}^{(k)}\right)\geq c\;\;\text{or}\;\;\mathrm{Ent}\left(\hat{\eta}_{l_{\mathrm{max}}}^{(k)}\mid\mu_{l}^{(k)}\right)=+\infty\}.$ (2.2)
@@
$l^{(k)}(c,l_{b})$ is the smallest level above $l_{b}$ such that, in the (unfortunate) case where the true target is given by the pessimistic choice $\check{\eta}_{l}^{(k)}$, a log cost of $c$ is necessary in order to perform importance sampling. For convenience, the dependency on $c$ and $l_{b}$ of $l^{(k)}$ will be removed if the context makes it clear. Note that the additional inequality constraint $\mathrm{Ent}\left(\hat{\eta}_{l_{\mathrm{max}}}^{(k)}\mid\mu_{l}^{(k)}\right)<+\infty$ in (2.2) ensures that the support of $\mu_{l}^{(k)}$ contains the support of the final target ${\eta}^{\ast}_{l_{\mathrm{max}}}$.
@@
In the present rare event case, the importance distribution is (1.3), the pointwise error bound is (1.2) and we propose the pessimistic and optimistic surrogate of the target (1.1)
@@
Using the surrogate (2.3), the particularization of the worst-case log cost in (2.2) is $\mathrm{Ent}\left(\check{\eta}_{l}^{(k)}\mid\mu_{l}^{(k)}\right)=\ln\frac{Z_{l}^{(k)}}{\check{Z}_{l}^{(k)}},$ thus the logarithm of the ratio of the approximate probability ${\pi\left(\left\{S^{(k)}>{l}\right\}\right)}$ over the smaller pessimistic probability ${\pi\left(\left\{S^{(k)}-E^{(k)}>{l}\right\}\right)}$.
@@
It can easily be checked that the critical level $l^{(k)}(c,l_{b})$ tends to infinity when the error $E^{(k)}$ vanishes pointwise.